OpenStudy (anonymous):

How do I integrate the following? int_{0}^{1}int_{0}^{x} e^{y/x}dydx I tried substituting the y/x part but didn't get very far :/

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\int\limits_{0}^{x} e^{y/x}dydx\]

OpenStudy (zarkon):

\[\int e^{y/x}dy=xe^{y/x}+c\]

OpenStudy (anonymous):

I managed to figure this out in the past while, but thanks anyway :)